Find the limit of {n^3/(n^4+1)} as n →∞


1
Expert's answer
2021-05-24T18:55:41-0400

"\displaystyle
\begin{aligned}
\lim_{n\to\infty} \frac{n^3}{n^4+1} &= \lim_{n\to\infty} \frac{1}{n + \frac{1}{n^3}} 
\\&= \lim_{n\to\infty} \frac{1}{n + 0} 
\\&= \lim_{n\to\infty} \frac{1}{n} = 0
\end{aligned}"